I am having a problem with simplifing complex fractions: for example: 3/4 + 1/8 over 5/12 - 1/6 can you explain?
Answer by edjones(8007)

(3/4 + 1/8) / (5/12 - 1/6)
=(6/8 + 1/8)/ (5/12 - 2/12)
= (7/8)/(3/12)
=(7/8)/(1/4)
=7/8 * 4
=7/2
.
Ed